OpenStudy (anonymous):

A right triangle has an area of 84ft^2 and a hypotenuse 25ft long. What are the lengths of the other two sides?

OpenStudy (mertsj):

\[\frac{1}{2}bh=84\] \[b^2+h^2=25^2\]

OpenStudy (mertsj):

Solve the system

OpenStudy (anonymous):

That's where I'm stuck at

OpenStudy (mertsj):

Solve the first equation for b and replace b in the second equation with that expression.

OpenStudy (anonymous):

I got 168/h^2+h^2=25^2

OpenStudy (mertsj):

Why is it 168/h^2?

OpenStudy (anonymous):

Because bh/2=84 so to get rid of the 2 I multiply 2 on both side and you get bh=168. To get b alone you divide h and both side then you said to plug it in

OpenStudy (mertsj):

So if you divide both sides by h, how do you get 168/h^2?

OpenStudy (anonymous):

Plug it in to b^2+h^2=25^2

OpenStudy (mertsj):

My point is that b = 168/h

OpenStudy (mertsj):

So, yes, plug it in.

OpenStudy (mertsj):

\[(\frac{168}{h})^2+h^2=625\]

OpenStudy (anonymous):

Okay then what?

OpenStudy (mertsj):

Square 168/h

OpenStudy (anonymous):

Okay then..

OpenStudy (mertsj):

1. Multiply both sides by h^2 2. Set the equation equal to 0 3. Factor if possible 4. If it won't factor, plug it into the quadratic formula.